RC Meeting and Run Plan, Wed, January 22nd 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

General updates

Target training on Thursday 1/23

  • Thursday from 2:30PM to 4:00PM in counting house.
  • Please send an email to JP (jpchen@jlab.org) or Arun (arunts@jlab.org) to register.

Tuesday day – Wednesday owl

  1. Laser alignment and wavelength adjustment
  2. New correction coil settings [1]
    • AFP loss per pass: 1.6% in pumping chamber, 0.2% in target chamber [2]
  3. NMR field sweep program's transverse pumping chamber signal works [3]
  4. pNMR work continues [4]

Wednesday day

  1. Multiple accelerator problems overnight and today; some beam this morning
  2. Harp scan around noon
    • IHA3H07A IHA3H07B YAG
    • x width at target ~550 um [5]; want <~350 um
    • This morning, MCC asked Simona if we can change beam position for tuning; acceptable procedure?
    • Changed positions back to nominal @12:45PM [6]
  3. Moller & spin dance postponed

Last few polarization measurements

spin up curve

  4% @  3AM Wed
 21% @  7AM Wed
 25% @  9AM Wed
 29% @ 12PM Wed 
 32% @  3PM Wed

Current Run Plan

  • Take NMR every 5–6 hours
  • Always take 120Hz of EDTM

Once beam is back

  1. Tune beam; want width < ~350 um, comparable to this entry
  2. Check beam positions [7]
  3. If during day: Moller & spin dance (pending Hall A status)
  4. Do NMR measurement
  5. Take data:
    if (3He polarization > 38%):
    polarized 3He production data; transverse data; 90degrees
    3He reference cell data

Wednesday swing – Thursday owl

  1. See above beam comments
  2. Track polarization to see if changing laser power helped
    • Take 2 more NMR measurements at 6PM and 9PM
    • After these 2 measurements, change polarization angle for transverse production data

Thursday day

  • DOE tour in the morning
    • 8:00 AM RadCon sweep
    • 9:30 AM restricted access
    • 10:30 AM tour, over by lunch presumably
  • Afternoon
    • See above beam comments


  • Arun tentatively getting replacement cell prepared early next week; install it in two(?) weeks?

Special Requests

  • Keep target expert schedule filled out on whiteboard

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[8]
  • Temporary Laser instructions can be found here [9]

Planned Accelerator Configuration Changes